OpenStudy (anonymous):

A rectangle is inscribed with its base on the x -axis and its upper corners on the parabola y=2-x^2 What are the dimensions of such a rectangle with the greatest possible area? width= height= i figured out height was 4/3 cause i thought width was sqrt(2/3) and plugged it in y=2-x^2

OpenStudy (amistre64):

Area = xy y = 2-x^2 Area = x(2-x^2) Area = 2x - x^3 derive and conquer :)

OpenStudy (amistre64):

A' = 2 - 3x^2 2 -3x^2 = 0 2 = 3x^2 2/3 = x^2 +-sqrt(2/3) = x

OpenStudy (amistre64):

x = sqrt(6)/3 y=2-(sqrt(6)/3)^2 y = 2 - (6/9) y = 18/9 - 6/9 y = 12/9 = 4/3 Max area: 4sqrt(6) ------ 9 i think :)